OpenStudy (exoexoexo):

if x-5, 2x, and 4x+6 are the first 3 terms of an arithmetic sequence, find x and the values of the three terms

satellite73 (satellite73):

what is \(2x-(x-5)\)?

OpenStudy (exoexoexo):

x-5

OpenStudy (exoexoexo):

wait x+5

satellite73 (satellite73):

yes

satellite73 (satellite73):

so the difference between the first two terms is \(x+5\)

satellite73 (satellite73):

now what is \(4x+6-2x\)?

OpenStudy (exoexoexo):

2x-6

satellite73 (satellite73):

i agree and since it is arithmetic, the differences must be the same to find \(x\) solve \[x+5=2x-6\]

OpenStudy (exoexoexo):

x=11

satellite73 (satellite73):

that is what i get too

satellite73 (satellite73):

now that you know \(x=11\) finding the value of the first 3 terms should be easy enough right?

OpenStudy (exoexoexo):

\[a _{1}=6\] \[a _{2}=22\] \[a _{3}=50\]

OpenStudy (exoexoexo):

that doesn't seem right

satellite73 (satellite73):

no it doesn't does it

OpenStudy (exoexoexo):

can we try \[2x=(x-5)+(2-1)(x+5)\]

OpenStudy (exoexoexo):

that doesn't work out either

satellite73 (satellite73):

hold on let me think see if i can figure out what is wrong with this method common differences must be equal

satellite73 (satellite73):

oh cause i am an idiot

satellite73 (satellite73):

i wrote \[x+5=2x-6\] but it really should have been \[x+5=2x+6\] doe

satellite73 (satellite73):

since \(4x+6-2x=2x+6\)

satellite73 (satellite73):

works now i bet

OpenStudy (exoexoexo):

x=-1

satellite73 (satellite73):

yes, now try it, see if it works

OpenStudy (exoexoexo):

thank you